A father is 30 years older than his son, one year ago he was 4 times as old as his son, find their present ages? Answer by mananth(16949) 
 son age =x
father -age = x+30
last years
son - age = x-1
father-age = (x+30)-1= x+29
(x+29)= 4(x-1)
x+29 = 4x-4
4x-x= 29+4
3x=33
/3
x= 11 years son's age
father's age = 30+11= 41 years